Find the equation of circle having (10,2) and (6,-4) as ends of diameter.
Answer by Cromlix(4381)

Hi there
First find the mid point of the
diameter. This will give you the
centre of the circle.
Mid point = x1 + x2/2, y1 + y2/2
Mid point = 10 + 6/2 , 2 + (-4)/2
Mid point = 16/2 , -2/2
Mid point = 8 , -1
Distance from (8,-1) to (10,2)
will give the radius
Distance formula = √(x2 - x1)^2 + (y2 - y1)^2
Distance formuls = √(10 - 8)^2 + (2-(-1))^2
Distance formula = √(2)^2 + (3)^2
Distance formula = √ 4 + 9
Distance formula = √ 13
Radius = √13
Circle formula:
(x - a)^2 + (y - b)^2 = r^2
Using Midpt (8, -1) and radius = √13
(x - 8)^2 + (y + 1)^2 = 13
Hope this helps :-)